OpenStudy (anonymous):

Given that cosθ = -11/61 and θ is in Quadrant III, match the following items. 1. sinθ 11/60 2. tanθ 60/11 3. cscθ -61/60 4. secθ -61/11 5. cotθ -60/61

OpenStudy (er.mohd.amir):

only tan is +ve in III Qua.

OpenStudy (anonymous):

your sure

OpenStudy (fedorable):

你他妈的所有IM出

OpenStudy (er.mohd.amir):

sin x=-60/61 tan x=60/11 cot x=11/60 cse x=-61/60 sec x=-61/11